OpenStudy (anonymous):

There are three kinds of cookies in a cookie jar: chocolate chip, peanut butter and oatmeal. If the number of oatmeal cookies is 2/3 the number of chocolate chip cookies and the number of chocolate chip cookies is 3/7 the number of peanut butter cookies, what fraction of the cookies are oatmeal?

OpenStudy (anonymous):

You want to set up your system of equations. Use \(c\), \(p\), and \(o\) as variables to represent the numbers of chocolate chip, peanut butter, and oatmeal cookies respectively. What algebraic expressions can you get?

OpenStudy (anonymous):

"the number of oatmeal cookies is 2/3 the number of chocolate chip cookies" expresses an equation for variables \(o\) and \(c\).

OpenStudy (anonymous):

"the number of chocolate chip cookies and the number of chocolate chip cookies is 3/7 the number of peanut butter cookies" this gives you another equation

OpenStudy (anonymous):

o=2/3c c=3/7p p=7/2o

OpenStudy (anonymous):

Now, just by common sense we know the total number of cookies is \(c + o + p\).

OpenStudy (anonymous):

You want to use the equations you wrote down to change the terms \(o\) and \(p\) into terms of \(c\).

OpenStudy (anonymous):

For example: \[ c + o + p = c+\left( \frac{2}{3} c \right) + p \]because \( \large o = \frac{2}{3} c\)

OpenStudy (anonymous):

@ArkGoLucky Do you understand?

OpenStudy (anonymous):

do I have to solve for all the variables?

OpenStudy (anonymous):

So: \[ c+o+p= 1 \]because the total cookies is 1 times the total cookies.

OpenStudy (anonymous):

No, you just need to get everything in terms of c

OpenStudy (anonymous):

wio, please forgive the interuption, but I saw your formula, and I suck at math... but i think i figured out the answer... do you mind if i stick around to see if im right?

OpenStudy (anonymous):

@Lurch No problem.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

So if I solve everything in terms of c, how does that help me find o

OpenStudy (anonymous):

oooh, I see what you mean.... nevermind, get it all in terms of \(o\).

OpenStudy (anonymous):

crap... I was wrong.... back to the scratch paper...

OpenStudy (anonymous):

and also, why does c+p+o=1 aren't o, c, p the number of cookies

OpenStudy (anonymous):

you're right... I was thinking a bit too far ahead...what I should have said is: \[ c+p+o = n \]Where \(n\) is the total number of cookies. What we want is to find \(o=??? n\). We use the fact that \(n = c+p+o\).

OpenStudy (anonymous):

is the answer 1/6

OpenStudy (anonymous):

@ArkGoLucky Just did the problem, and that is what I got.

OpenStudy (anonymous):

okay thanks for the help

OpenStudy (anonymous):

@ArkGoLucky http://www.wolframalpha.com/input/?i=o+%3D+%282%2F3%29c%2C+c+%3D+%283%2F7%29p%2C+c+%2B+p+%2B+o+%3D+n%2C+b+%3D+o%2Fn Here is proof that you are correct. I make another variable \(b\) which was equal to \(o/n\). Wolfram said \(b = 1/6\).

OpenStudy (anonymous):

thank you. It feels good to be right